The 95% lower confidence bound for a population parameter θ is 9.232. The point estimate from the sample used to construct this bound is 15.7. Use this information to compute the standard error in the sample in both of the following cases.

a) If this is a two-sided confidence interval, then the standard error is:

b) If this is a one-sided confidence interval, then the standard error is:
Answer #1

Ans:

b)

Margin of error=15.7-9.232=6.468

one sided z value for 95% confidence level=1.645

standard error=6.468/1.645=3.932

a)

two sided z value=1.96

standard error=6.468/1.96=3.300